Hayhayz (hayhayz):

Angie is working on solving the exponential equation 23x = 6; however, she is not quite sure where to start. Using complete sentences, describe to Angie how to solve this equation. @imqwerty

Hayhayz (hayhayz):

\[23^x \]

imqwerty (imqwerty):

use this rule- If \(a^x=b\) then \(log_ab=x\)

Hayhayz (hayhayz):

http://prntscr.com/af8auu ?

imqwerty (imqwerty):

yes thats correct :)

Hayhayz (hayhayz):

Theres not really much to describe in complete sentences though ey? lol thanks last questionn

imqwerty (imqwerty):

okay we can tell her to do this- since the given equation is in exponential form it is better to remove the exponential part to solve the equation and to do this we must take logarithm on both sides \(log(23^x)=log(6)\) then use this identity- \(\color{red}{log(a^b)=blog(a)}\) xlog(23)=log(6) and then we can ask her to simplify the rest :) x=\(\large \frac{log(6)}{log(23)}\) x=\(log_{23}(6)\)

Hayhayz (hayhayz):

Thanks! :D

imqwerty (imqwerty):

np :)